Shift report 20030514 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed May 14 2003 - 16:11:49 EDT

  • Next message: Flemming Videbaek: "short term schedule update"
    ------------------------------------------------------------
    Date: Wed May 14 16:11:49 2003, Shift: 08:00-16:00
    Supervisor: fv, with: 
    Log book#14 pages 101 - 101
    
    Shift Summary:
    ------------------------------------------------------------
    1) Beam from 8-9.15
       Again from ~ 13:15 to 16:00
    
       Though not so useful since many magnets problems.
    
       D2 tripped just while starting and did not notice.
    
       D1 had a bourning foul smeel. Magnet guys repaired this from
       11.00 to 15:15
    
    2) Useful runs only 8958 1/1A 4 deg started at 15:40.
       Additional timing of spin scalers.
    3) Si still off
    
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed May 14 2003 - 16:12:46 EDT